OpenStudy (anonymous):

A tree grows 1 3/4 feet per year. How long will it take the tree to grow from a height of 21 1/4 feet to a height of 37 feet?

OpenStudy (anonymous):

wow we gotta do a bunch of arithmetic here

OpenStudy (anonymous):

first we need how much it grows' what is \[37-21\tfrac{1}{4}\]?

OpenStudy (anonymous):

I'm asking my friend and she is telling me a lot of complicated steps, and she says the answer is some how 14?

OpenStudy (anonymous):

it takes two steps, although the division will take a couple in itself

OpenStudy (anonymous):

Do I have to find a common denominator?

OpenStudy (anonymous):

you have to subtract is all

OpenStudy (anonymous):

as a first step

OpenStudy (anonymous):

you can probably compute \[37-21\tfrac{1}{4}\] in your head

OpenStudy (anonymous):

like if you have thirty seven dollars and you spend twenty one and a quarter, how much money do you have left?

OpenStudy (anonymous):

that equals 16

OpenStudy (anonymous):

you still have to subtract a quarter

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

you have $37 you spend $21.25 \[37-21=16\] but \[1-\frac{1}{4}=\frac{3}{4}\] so you have only \[\$15.75\] left in terms of money with fractions you have \[15\tfrac{3}{4}\]

OpenStudy (anonymous):

now you have to compute \[15\tfrac{3}{4}\div1\tfrac{3}{4}\]

OpenStudy (anonymous):

I have 8 whole and 4/7